The sum of the present age of Vatha and Chris is 36. In 4 years time, the sum of their ages will equal twice Vatha's persent age. How old are they now?
Answer by ad_alta(240)

Let 'V' be Vatha's age and 'C' Chris' age. Then V+C=36 and (V+4)+(C+4)=2*V. Substituting the first equation into the second, we get (V+4)+(36-V+4)=2*V, so V=22 and C=14.